suppose a math class contains 25 students, 14 females (three of whom speak French) and 11 males (two of whom speak French. Compute the probability that a randomly selected student is male, given that the student speaks French.
Answer #1

Answer:

To determine the probability that a randomly selected student is male

Let us assume M for male,F for female & R is french

Now consider,

P(M) = 11/25

= 0.44

P(F) = 14/25

= 0.56

P(R/M) = 2/11

= 0.1818

P(R/F) = 3/14

= 0.2143

Now let us consider the baye's theorem,

P(M/R) = P(M)*P(R/M) /(P(M)P(R/M) + P(F)P(R/F))

now substitute the given known values in above expression

P(M/R) = 0.44*0.1818 / (0.44*0.1818 + 0.56*0.2143)

P(M/R) = 0.4